OpenStudy (anonymous):

HELP!! A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days. f(n) = 12(1.03)^n What does the y-intercept of the graph of the function f(n) represent?

OpenStudy (anonymous):

@SolomonZelman

jimthompson5910 (jim_thompson5910):

Think of a visual graph. Where does the y-intercept happen on this graph?

OpenStudy (anonymous):

on the y axis

jimthompson5910 (jim_thompson5910):

what is x equal to for any point on the y axis?

OpenStudy (anonymous):

0

jimthompson5910 (jim_thompson5910):

so in this case, they're using n instead of x (doesn't matter really)

jimthompson5910 (jim_thompson5910):

so the y intercept happens when n = 0

jimthompson5910 (jim_thompson5910):

plug in n = 0 into the function and tell me what you get

OpenStudy (anonymous):

y=12

jimthompson5910 (jim_thompson5910):

so the y-intercept is the point (0,12) meaning that on day 0, ie the starting day, the height of the plant is 12 cm

jimthompson5910 (jim_thompson5910):

basically the starting height is 12 cm

OpenStudy (anonymous):

okay thank you!!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

The y intercept of the graph of the function f(n) represents that when n=0 the height is 12. So when there are 0 days the height of the plant will be 12 cm. is that okay?

jimthompson5910 (jim_thompson5910):

The last sentence is a bit odd. I'd say, on day 0, the height is 12 cm.

jimthompson5910 (jim_thompson5910):

But you have it down for the most part.

OpenStudy (anonymous):

oh okay, can you help me with one more corresponding question?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

What is the average rate of change of the function f(n) from n = 3 to n = 10 and what does it represent?

jimthompson5910 (jim_thompson5910):

The average rate of change is simply the slope of the line through the two points where n = 3 and n = 10

jimthompson5910 (jim_thompson5910):

So you first need to find f(3) and f(10)

OpenStudy (anonymous):

kay, one moment please

OpenStudy (anonymous):

f(3)= 13.11 and f(10)= 16.13

jimthompson5910 (jim_thompson5910):

Those are basically what I'm getting as well (just with longer decimals)

jimthompson5910 (jim_thompson5910):

f(3) = 13.11 means you have the point (3, 13.11) f(10) = 16.13 means you have the point (10,16.13)

jimthompson5910 (jim_thompson5910):

You have these two points (3, 13.11) & (10,16.13) Finding the slope through them is the same as finding the avg rate of change from n = 3 to n = 10

OpenStudy (anonymous):

okay so i'll do 16.13-13.11/10-3?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

I get 0.431428714

jimthompson5910 (jim_thompson5910):

I'm getting the same if you compute (16.13-13.11)/(10-3) It turns out that a more accurate answer is 0.43061036458991 So you're correct to 2 decimal places

jimthompson5910 (jim_thompson5910):

To get 0.43061036458991, I computed (16.1269965521294-13.112724)/(10-3)

OpenStudy (anonymous):

yes that is what I got can you help me with another question if you want?

jimthompson5910 (jim_thompson5910):

go for it

OpenStudy (anonymous):

The price of products may increase due to inflation and decrease due to depreciation. Derek is studying the change in the price of two products, A and B, over time. The price f(x), in dollars, of product A after x years is represented by the function below. f(x) = 0.69(1.03)^x Part A: Is the price of product A increasing or decreasing and by what percentage per year? Justify your answer.

jimthompson5910 (jim_thompson5910):

Ok post any thoughts or how far you got

OpenStudy (anonymous):

okay I plugged in numbers 1,2,3,4 for the exponent

jimthompson5910 (jim_thompson5910):

and what did you notice?

OpenStudy (anonymous):

then gathered the info of f(1)= 0.7107, f(2)=0.732021, f(3)= 0.75398163 , f(4)=0.7766010789

OpenStudy (anonymous):

that's where I stopped

jimthompson5910 (jim_thompson5910):

So is the price of product A increasing over time? or decreasing over time?

OpenStudy (anonymous):

increasing but it asks for percentages

jimthompson5910 (jim_thompson5910):

hint: try to write 1.03 in the form 1+r

jimthompson5910 (jim_thompson5910):

r will be the rate of increase

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

what do i multiply or divide?

jimthompson5910 (jim_thompson5910):

we can break up 1.03 into 1+0.03 that matches the form 1+r where r = 0.03

jimthompson5910 (jim_thompson5910):

So product A is increasing in price 3% each year (ie inflation for this product is 3%)

OpenStudy (anonymous):

you divided to find the percentage?

jimthompson5910 (jim_thompson5910):

no I just broke up 1.03 into 1+0.03 Then you take 0.03 and multiply it by 100 to get 0.03*100 = 3%

OpenStudy (anonymous):

okay, how will i find the percentage for part B?

jimthompson5910 (jim_thompson5910):

find the difference in price from year 1 to year 2

jimthompson5910 (jim_thompson5910):

then divide that by the price in year 1 finally, multiply that result by 100 to get a percentage

OpenStudy (anonymous):

so it will be10,201-10,100= 101

OpenStudy (anonymous):

then 101/10100= 0.01

jimthompson5910 (jim_thompson5910):

so far, so good

OpenStudy (anonymous):

0.01*100= 1%

jimthompson5910 (jim_thompson5910):

So the price increased by 1% from year 1 to year 2 (for product B)

jimthompson5910 (jim_thompson5910):

What's the percentage increase from year 2 to year 3?

OpenStudy (anonymous):

okay i get 1%

jimthompson5910 (jim_thompson5910):

how about from year 3 to year 4?

OpenStudy (anonymous):

1%

OpenStudy (anonymous):

well i got 0.999999 but i rounded it

jimthompson5910 (jim_thompson5910):

so basically the inflation is 1% throughout each year

OpenStudy (anonymous):

yes thank you

jimthompson5910 (jim_thompson5910):

np